Better Mousetraps has come out with an improved product, and the world is beating a path to its door. As a result, the firm projects growth of 21 percent per year for 4 years. By then, other firms will have copycat technology, competition will drive down profit margins, and the sustainable growth rate will fall to 6 percent. The most recent annual dividend was DIV0 = $2.9 per share. What is the stock price today if the discount rate is 9 percent?
Answer #1

D1=(2.9*1.21)=3.509

D2=(3.509*1.21)=4.24589

D3=(4.24589*1.21)=5.1375269

D4=(5.1375269*1.21)=6.21640755

Value after year 4=(D4*Growth rate)/(Discount rate-Growth rate)

=(6.21640755*1.06)/(0.09-0.06)

=219.6464

Hence stock price=Future dividend and value*Present value of discounting factor(rate%,time period)

=3.509/1.09+4.24589/1.09^2+5.1375269/1.09^3+6.21640755/1.09^4+219.6464/1.09^4

=$170.77(Approx).